Newton's Laws, Work & Kinetic Energy

You weigh 530 N. Your bathroom scale contains a light but very stiff ideal spring. When you stand at rest on the scale, the spring is compressed 1.80 cm. Your 180 N dog then gently jumps into your arms.


How much work is done by the spring as the two of you are brought to rest by friction?

Express the answer to three significant figures and include the appropriate units.
Answer #1

SOLUTION :


Spring constant 

= my weight / compression

= 530 / 1.80 kg/cm

= 294.44 kg/cm


By putting additional weight pf dog of 180 N, additional compression of spring

= 180/294.44

= 0.61132 cm 

= 0.061132 m 


So,


Work done 

= Total weight of mine and dog moves by additional compression

= (530 + 180) * 0.061132 N-m 

= 43.404 N-m

= 43.404 J (ANSWER).

answered by: Tulsiram Garg